# First Dark Photon Search Results from the Dandelion Experiment

### v1

@@ Abstract
This paper presents the first results from the Dandelion experiment, a directional detection, which searches for 1 meV dark photon dark matter. We use a spherical mirror to convert dark photons into standard millimeter-wavelength photons that can then be detected with an array of 221 Kinetic Inductance Detectors (KIDs) cooled down to 150 mK within the KISS (KIDs Interferometric Spectral Surveyor) camera and operating between 150 and 350 GHz. We used 1480 minutes of data to search for the signal of dark photons in the KID detectors, which is expected to be modulated due to the Earth’s rotation. Our main challenge was to deal with a large background from room temperature and stray-light fluctuations. We used a de-correlation analysis to remove these background fluctuations. Templates of the background fluctuations were constructed from a Principal Component Analysis decomposition of detector measurements outside the expected Field of View trajectory of dark photons. We found that the dark photon signal was consistent with zero, giving a new upper limit on the dark photon’s kinetic mixing, $\chi$, with masses between 0.6 meV and 1.4 meV. These are the first constraints on dark photons as a dark matter candidate using an array of KIDs at millimeter wavelength.
@@ 1 Introduction
Different scenarios depending on whether the symmetry is broken or not during inflation were considered in [20]. If the symmetry is already broken during inflation and the DH is heavy, i.e. $H_{I}<m_{h}$, $H_{I}$ being the inflationary Hubble parameter, the main contribution to the relic DM abundance comes anyway from inflationary production, with that from the Higgs decay being subdominant. Whereas in the case of a light DH, $m_{h}<H_{I}$, one can have, for example, resonant production of the DP due to the oscillations of the DH when inflation ends [9]; however this would be suppressed for an meV DP that will privilege the scenario when the symmetry is broken after inflation and $H_{I}>v$. The production can be related to the dynamics of the dark Higgs [21, 20], or to the cosmic strings associated to the breaking of symmetry [18, 15, 16]. Cosmic strings will also give rise to a stochastic background of gravitational waves, which could be detectable by future gravity waves detectors.
@@
The main difficulty is finding a very weak signal that is hidden in a large amount of background noise. This noise affects all our pixels simultaneously. To solve this, we use a method called Principal Component Analysis (PCA). We apply PCA to the data from the "Background" pixels to learn the main patterns of the noise. This allows us to build a precise noise model and remove it from all our data, making it possible to search for the signal underneath. A key advantage of our strategy is the rapid modulation between two mirror positions every second. This ensures that pixels capturing the signal in one position provide a direct measurement of the background noise in the alternative position just a second later. This quasi-simultaneous reference allows us to monitor the background of the signal pixels at all times, providing the ’clean’ data necessary for an accurate PCA-based noise decomposition.
@@ 2 Experimental Setup
In this experiment, we use a spherical aluminum mirror with a diameter of 50 cm and a radius of curvature of 5 meters. This mirror serves as the conversion surface. If a dark photon from the galactic halo hits the mirror, it can be converted into a standard photon with an energy equal to the dark photon’s mass, corresponding to a signal with a predicted wavelength of 1.24 mm. These newly created photons are emitted nearly perpendicular to the mirror’s surface. To detect those photons, we use an array of 221 Kinetic Inductance Detectors (KIDs) [24], each with a sensitive surface area of 2.8×2.8 $\times 2.8\text{ mm}^{2}$. These are superconducting microwave resonators that change their resonant frequency when they absorb energy, allowing the energy of an incident photon to be measured. The array is cooled to 150 mK inside a cryostat to minimize thermal noise (Fig. 1). The system is optimized for frequencies around 250 GHz (corresponding to a wavelength of $\lambda\approx 1.20$ mm and a DP mass of $m_{\chi}\approx 1.03$ meV), aligning with the expected signature of 1 meV dark photon dark matter. The detector bandwidth ranges from 150 GHz ($\lambda\approx 2.00$ $m$, $m_{\chi}\approx 0.62$ meV) to 350 GHz ($\lambda\approx 0.86$ $m$, $m_{\chi}\approx 1.45$ meV); see Appendix A.
@@
Figure 1: The figure shows a schematic diagram of the experimental setup [6]
@@
During the 1480-minute measurement, we collected 148 data files of 10 minutes each. The mirror was switched between two positions, a "frontal" and a "tilted" alignment. This change was made every second within each 10-minute data file. This rapid modulation technique is designed to mitigate low-frequency instrumental drifts, such as 1/f noise. For the analysis, we use a mask to separate the data into two distinct time series, one for each mirror position. For the analysis, we concatenate the data from both mirror positions into a single, unified time series. This approach allows us to perform a global analysis on the entire 1480-minute dataset, effectively doubling the integration time compared to analyzing positions separately.
@@ 3 Signal modeling and reconstruction procedure
For a given mirror position, we measure a change in the resonance frequency of each pixel in the KID array (this frequency shift $\Delta f$ corresponds to the energy of the absorbed photon). These frequency measurements (in Hz) for each pixel are then converted into temperature by multiplying by the pixel’s response (calibrated in K/Hz). The resulting time series for each pixel shows the evolution of the equivalent temperature shift corresponding to the power absorbed by that pixel during the observation period (1480 minutes on January 11, 2024). We have 148 data files, each lasting 10 minutes. To study the evolution of the absorbed temperature (or frequency) without being affected by cosmic glitches, we take the median over each 10-minute file, giving one point per file for a chosen position. The absorbed power comes from several background sources (room temperature, thermal emission from the mirror and optical elements, stray light, etc.), in addition to the possible desired signal: standard photons produced from the conversion of dark photons.
@@ 3.1 Expected Dark Photon Signal
$\rho_{\rm CDM}$ is the local dark matter density, $\eta(E_{\gamma})$ is the detection efficiency, $A_{\rm mirr}$ the surface area of the mirror, $I(x,y,t)$ the normalized signal spread on the detector, and $\alpha(t)$ the angle between the dark photon polarization and the tangent of the mirror, we use $\langle\cos^{2}\alpha(t)\rangle_{T}\approx\frac{2}{3}$ [6]. For the purpose of our analysis, it is useful to separate this equation into two parts: an overall power amplitude and the spatial distribution $I(x,y,t)$. We can define the power amplitude, as:
@@
While the exact mathematical form of a Fresnel diffraction pattern is complex, its central lobe—where most of the signal power is concentrated—can be accurately approximated by a two-dimensional Gaussian function. To precisely define our signal model, we performed a fit of the theoretical Fresnel diffraction profile using a normalized two-dimensional Gaussian function, as illustrated in Fig. 3. The fit gives a standard deviation of $\sigma_{D}=3.2\pm 0.04\penalty 10000\ mm$. For the purposes of our analysis, we adopt this empirically determined value as a fixed parameter for the signal width (diffraction). Modeling the signal profile as a Gaussian offers two key advantages: it is analytically straightforward and computationally efficient for the final fitting procedure.
@@ 3.2 Background contribution
The evolution of temperature (Fig. 4) shows a correlation between pixels caused by uniform thermal background noise that affects the entire array. The main contribution arises from the thermal emission of the mirror and the surrounding structure at room temperature, which continuously radiates in the millimeter range. Additional power is emitted and reflected by the optical lenses located at different temperature stages (300 K, 4 K, and 150 mK). Even though the coldest lenses emit very little, the warmer ones still send a noticeable flux toward the detectors. A smaller contribution comes from electronic and readout noise, as well as residual stray light entering the cryostat through imperfect shielding or reflections. All these sources vary slowly with time because of temperature changes in the cryostat, or fluctuations in the electronics, and because the pixels see the same optical Field of View (FOV), their signals show a strong temporal correlation. This fact produces temperature fluctuations of the order of kelvins, completely hiding the expected dark photon signal.
Figure 4: The left and middle panels show the raw temperature time series for a subset of ’Signal’ pixels and the full set of ’Background’ pixels, respectively. Both are dominated by large, correlated thermal fluctuations. To model this common-mode noise, we apply Principal Component Analysis (PCA) to the concatenated background data. The right panel displays the first five (out of 10 used) resulting principal components, which represent the dominant, independent modes of temporal variation. These components are used to construct the noise model for subtraction
@@ 3.3 Two-region model
To isolate the potential dark photon signal from the dominant background noise, we construct a model for the temperature measured by each pixel over time. Our model assumes that the data from any given pixel is a sum of a background component and a potential dark photon induced signal component. The background component, which affects all pixels, is described as a linear combination of time-dependent noise templates, $T^{\mathrm{syst}}_{i}(t)$:
@@
where $T_{S}$ is the dark photon associated signal amplitude in Kelvin, $(x(t),y(t))$ is the signal trajectory, and $(x_{\text{k}},y_{\text{k}})$ represents the position of the pixel in the array. The coefficients $b_{k}$ (offset) and $a_{k,i}$ (slopes) are unique to each pixel "k" and define its specific response to these shared background fluctuations. The signal component, $S(x_{k}(t),y_{k}(t))$, is different. We only expect it to appear in the pixels that fall along the predicted signal path at a given time t.
@@
To solve this problem, we apply Principal Component Analysis (PCA) [22] to the pixel data from the “Background” region for each mirror’s position. PCA is a linear decomposition that identifies the common temporal variation modes (the principal components, $T^{\mathrm{syst}}_{i}(t)$) which capture the majority of the correlated background noise. Modeling the background noise using PCA is a powerful technique to extract patterns of common variation (Fig. 4) from a complex, correlated dataset. In our case, the data consist of the temporal time series of pixels in the “Background” region, where we are confident that only noise is present.
The robustness of our PCA templates relies on the two-position modulation technique. By switching the mirror alignment every second, we effectively observe the background for every pixel during its ’off-signal’ intervals. Consequently, for the pixels located on the signal trajectory, the data collected in the alternate mirror position serves as a high-fidelity reference of the environmental noise. These background-only intervals from all pixels are concatenated to train the PCA, ensuring the noise model is perfectly representative of the conditions during signal acquisition.
We select 190 pixels in the background region. We concatenate their measurements from both mirror positions to form a single dataset and decompose the signal into independent components using a Principal Component Analysis (PCA). Notice that these independent components will include both correlated contributions across all detectors and the noisy contributions. Therefore, the choice of the number of Principal Components ($N_{\text{sys}}$) used to model the background noise (correlated between detectors) is a critical step in the analysis. This choice is guided by the goal of minimizing the final uncertainty on the signal amplitude ($\sigma_{T}$). Although the noise characteristics can differ slightly between the two datasets, leading to slightly different optimal PC counts, we chose a single number of components for consistency. As shown in Figure 6, we selected $N_{\text{sys}}=10$ for both mirror positions. This value represents a robust choice, located in the region where the signal uncertainty is at its minimum, ensuring a balanced and effective noise subtraction.
@@ 3.4 Final analysis
To fit our model to the data, we use a maximum likelihood estimation (MLE) framework. For each pixel, we construct a likelihood function based on the assumption of Gaussian noise.
By maximizing this function, we find the best-fit values for all the model parameters: the offset, the contribution of each Principal Component, and the amplitude of the signal. This provides a robust noise model for each pixel, which can then be subtracted to search for a significant signal. For pixels near the expected trajectory, the signal component, proportional to the Gaussian in Eq.(6), is included in the fit, allowing us to directly constrain its amplitude.
@@
Figure 7: Performance of the noise subtraction procedure. Left: Time evolution of the raw temperature data (solid line) compared to the model (dashed line) for a representative pixel, showing the model’s ability to accurately track large-scale thermal fluctuations. Center: Resulting temperature residuals for all pixels in the signal region after model subtraction; the fluctuations are centered around zero with a significantly reduced amplitude compared to the raw data. Right: Statistical distribution of the combined residuals across all pixels, fitted with a Gaussian profile to demonstrate that the remaining noise is stochastic.
@@ 4 Exclusion Curve
The next step is to translate this temperature limit into a limit on the power incident on the detector. The relationship between power and temperature variation $\Delta T$, for a blackbody detector, is described by Stefan-Boltzmann’s law, integrated over the bandwidth $H(\nu)$, $A\Omega$ is the throughput of each pixel, $\zeta(\nu)$ is the efficiency of the optical system composed of the three lenses $\langle\zeta(\nu)\rangle=0.6$, and $\langle\eta(\nu)\rangle=\langle\zeta(\nu)\cdot H(\nu)\rangle=0.3$ represents the detection efficiency seen in equation 4. The power amplitude is given by:
$A_{\text{power}}=A\Omega\int_{\nu_{min}}^{\nu_{max}}\zeta(\nu)H(\nu)B(\Delta T,\nu)d\nu.$ (7)
In our operating regime, where the total system temperature is high relative to the photon energy we want to measure ($k_{B}T_{\text{tot}}\gg h\nu$), we can use the Rayleigh-Jeans approximation, so we have $B(\Delta T,\nu)=B(T_{background}+\Delta T,\nu)-B(T_{background},\nu)$. This condition is largely satisfied, especially for the 300 K lens, where the thermal energy is around 25 times greater than the energy of the 244 GHz photon. By calculating the integral, taking into account the measured bandwidth and the limit $\Delta T_{\text{lim}}$, we can convert the temperature limit into a detectable power limit. This limit on power is then converted into a limit on kinetic mixing coupling $\chi$. The final expression for the coupling is:

### v2

@@ Abstract
This paper presents the first results from the Dandelion experiment, a directional detection experiment, which searches for 1 meV dark photon dark matter. We use a spherical mirror to convert dark photons into standard millimeter-wavelength photons that can then be detected with an array of 221 Kinetic Inductance Detectors (KIDs) cooled down to 150 mK within the KISS-NIKA camera (KIDs Interferometric Spectral Surveyor) and operating between 150 and 350 GHz. We used 1480 minutes of data to search for the signal of dark photons in the KID detectors, which is expected to be modulated due to the Earth’s rotation. Our main challenge was to deal with a large background from room temperature and stray-light fluctuations. We used a de-correlation analysis to remove these background fluctuations. Templates of the background fluctuations were constructed from a Principal Component Analysis decomposition of detector measurements outside the expected Field of View trajectory of dark photons. We found that the dark photon signal was consistent with zero, giving a new upper limit on the dark photon’s kinetic mixing, $\chi$, with masses between 0.6 meV and 1.4 meV. These are the first constraints on dark photons as a dark matter candidate using an array of KIDs at millimeter wavelength.
@@ 1 Introduction
Different scenarios depending on whether the symmetry is broken or not during inflation were considered in [27]. If the symmetry is already broken during inflation and the DH is heavy, i.e. $H_{I}<m_{h}$, $H_{I}$ being the inflationary Hubble parameter, the main contribution to the relic DM abundance comes anyway from inflationary production, with that from the Higgs decay being subdominant. Whereas in the case of a light DH, $m_{h}<H_{I}$, one can have, for example, resonant production of the DP due to the oscillations of the DH when inflation ends [13]; however, this would be suppressed for an meV DP that will privilege the scenario when the symmetry is broken after inflation and $H_{I}>v$. The production can be related to the dynamics of the dark Higgs [28, 27], or to the cosmic strings associated to the breaking of symmetry [23, 20, 21]. Cosmic strings will also give rise to a stochastic background of gravitational waves, which could be detectable by future gravity waves detectors.
@@
The main difficulty is finding a very weak signal that is hidden in a large amount of background noise. This noise affects all our pixels simultaneously. We can take advantage of the signal and background noise properties to separate them. As discussed above the DP signal will be modulated and as so it will vary in shape and amplitude with time and from pixel to pixel. By contrast the background noise varies with time but it is expected to be correlated within pixels.
@@ 2 Experimental Setup
In this experiment, we use a spherical aluminum mirror with a diameter of 50 cm and a radius of curvature of 5 meters. This mirror serves as the conversion surface. If a dark photon from the galactic halo hits the mirror, it can be converted into a standard photon with an energy equal to the dark photon’s mass, corresponding to a signal with a predicted wavelength of 1.24 mm. These newly created photons are emitted nearly perpendicular to the mirror’s surface. To detect those photons, we use an array of 221 Kinetic Inductance Detectors (KIDs) [31], each with a sensitive surface area of 2.8×2.8 $\times 2.8\text{ mm}^{2}$. KIDs are high-quality microwave resonators that have a change in their electromagnetic properties in response to incoming radiation. The resonator is a superconductor which absorbs photons exceeding its gap energy, resulting in Cooper pairs breaking and thus, a change in the ratio of paired and unpaired (quasi-particles) charge carriers, which changes their resonant frequency, allowing the energy of an incident photon to be measured. The array is cooled to 150 mK inside the KISS-NIKA cryostat [25, 24] to minimize thermal noise (Fig. 1). The system is optimized for frequencies around 250 GHz (corresponding to a wavelength of $\lambda\approx 1.20$ mm and a DP mass of $m_{\chi}\approx 1.03$ meV), aligning with the expected signature of 1 meV dark photon dark matter. The detector bandwidth ranges from 150 GHz ($\lambda\approx 2.00$ $m$, $m_{\chi}\approx 0.62$ meV) to 350 GHz ($\lambda\approx 0.86$ $m$, $m_{\chi}\approx 1.45$ meV); see Appendix A.
@@
Figure 1: Schematic diagram of the experimental setup [6], illustrating the operational DP-to-photon conversion process where a 5m-radius mirror emits signals toward the KISS-NIKA camera.
@@
During the 1480-minute measurement, we collected 148 data files of 10 minutes each. The mirror was switched between two positions, a "frontal" and a "tilted" alignment. This change was made every second within each 10-minute data file. For the analysis, we use a mask to separate the data into two distinct time series, one for each mirror position. For the analysis, we concatenate the data from both mirror positions into a single, unified time series. This approach allows us to perform a global analysis on the entire 1480-minute dataset, effectively doubling the integration time compared to analyzing positions separately.
@@ 3 Signal modeling and reconstruction procedure
For a given mirror position, using the NIKEL readout electronics [8, 9], we measure the change in resonance frequency $\Delta f$ for each pixel, which is proportional to the input energy of the absorbed photon [7, 16, 11].
These frequency measurements (in Hz) for each pixel are then converted into temperature by multiplying by the pixel’s response (calibrated in K/Hz). The resulting time series for each pixel shows the evolution of the equivalent temperature shift corresponding to the power absorbed by that pixel during the observation period (1480 minutes on January 11, 2024). We have 148 data files, each lasting 10 minutes. To study the evolution of the absorbed temperature (or frequency) without being affected by cosmic glitches, we take the median over each 10-minute file, giving one point per file for a chosen position. The absorbed power comes from several background sources (room temperature, thermal emission from the mirror and optical elements, stray light, etc.), in addition to the possible desired signal: standard photons produced from the conversion of dark photons.
@@ 3.1 Expected Dark Photon Signal
$\rho_{\rm CDM}$ is the local dark matter density, $\eta(E_{\gamma})$ is the detection efficiency, $A_{\rm mirr}$ is the surface area of the mirror, $I(x,y,t)$ the normalized signal spread on the detector, and $\alpha(t)$ the angle between the dark photon polarization and the tangent of the mirror, we use $\langle\cos^{2}\alpha(t)\rangle_{T}\approx\frac{2}{3}$ [6]. For the purpose of our analysis, it is useful to separate this equation into two parts: an overall power amplitude and the spatial distribution $I(x,y,t)$. We can define the power amplitude, as:
@@
While the exact mathematical form of a Fresnel diffraction pattern is complex, its central lobe—where most of the signal power is concentrated—can be accurately approximated by a two-dimensional Gaussian function. To precisely define our signal model, we performed a fit of the theoretical Fresnel diffraction profile using a normalized two-dimensional Gaussian function, as illustrated in Fig. 3. The fit gives a standard deviation of $\sigma_{D}=3.2\pm 0.04\penalty 10000\ mm$. For the purposes of our analysis, we adopt this empirically determined value as a fixed parameter for the signal width (diffraction). Modeling the signal profile as a Gaussian function offers two key advantages: it is analytically straightforward and computationally efficient for the final fitting procedure.
@@ 3.2 Background contribution
The evolution of temperature (Fig. 4) shows a correlation between pixels caused by uniform thermal background noise that affects the entire array. The main contribution arises from the thermal emission of the mirror and the surrounding structure at room temperature, which continuously radiate in the millimeter range. Additional power is emitted and reflected by the optical lenses located at different temperature stages (300 K, 4 K, and 150 mK). Even though the coldest lenses emit very little, the warmer ones still send a noticeable flux toward the detectors. A smaller contribution comes from electronic and readout noise, as well as residual stray light entering the cryostat through imperfect shielding or reflections. All these sources vary slowly with time because of temperature changes in the cryostat, or fluctuations in the electronics, and because the pixels see the same optical Field of View (FOV), their signals show a strong temporal correlation. And the most important fluctuations come from the mirror temperature and can lead to variations of the order of a Kelvin.
Figure 4: The left and middle panels show the raw temperature time series for a subset of ’Signal’ pixels and the full set of ’Background’ pixels, respectively. Both are dominated by large, correlated thermal fluctuations. To model this common-mode noise, we apply Principal Component Analysis (PCA) to the concatenated background data. The right panel displays the first five (out of 10 used) resulting principal components, which represent the dominant, independent modes of temporal variation. Note that these principal components have been normalized for visualization purposes, which is why their vertical scale does not visually correspond to the large, unscaled fluctuations seen in the raw data of the left and middle panels. These components are used to construct the noise model for subtraction.
@@ 3.3 Two-region model
To isolate the potential dark photon signal from the dominant background noise, we construct a model for the temperature measured by each pixel over time. Our model assumes that the data from any given pixel is a sum of a background component and a potential dark photon induced signal component (Eq. 6).
@@
The background component, which affects all pixels, is described as a linear combination of time-dependent noise templates, $T^{\mathrm{syst}}_{i}(t)$, $T_{S}$ is the dark photon associated signal amplitude in Kelvin, $(x(t),y(t))$ is the signal trajectory, and $(x_{\text{k}},y_{\text{k}})$ represents the position of the pixel in the array. The coefficients $b_{k}$ (offset) and $a_{k,i}$ (slopes) are unique to each pixel "k" and define its specific response to these shared background fluctuations. The signal component, $S(x_{k}(t),y_{k}(t))$, is different. We only expect it to appear in the pixels that fall along the predicted signal path at a given time t.
@@
To solve this problem, we apply Principal Component Analysis (PCA) [29] to the pixel data from the “Background” region for each mirror position. PCA is a decomposition that identifies the common temporal variation modes (the principal components, $T^{\mathrm{syst}}_{i}(t)$) which capture the majority of the correlated background noise. Modeling the background noise using PCA is a powerful technique to extract patterns of common variation (Fig. 4) from a complex, correlated dataset. In our case, the data consist of the temporal time series of pixels in the “Background” region, where we are confident that only noise is present. To ensure the stability of our analysis, we verified that performing two parallel and independent PCA analyses yielded equivalent background models and final exclusion limits across both mirror positions. While the two-position modulation allows us to observe background fluctuations every second, the decision to use a unified PCA approach was made to effectively capture common-mode noise that was otherwise difficult to decouple between the two individual positions. By concatenating these background-only intervals to train the PCA, we ensure that the resulting noise model is fully representative of the conditions during the entire acquisition period.
We select 190 pixels in the background-only region. We concatenate their measurements from both mirror positions to form a single dataset and decompose the signal into independent components using a Principal Component Analysis (PCA). Note that these independent components will include both correlated contributions across all detectors and the noisy contributions. Therefore, the choice of the number of Principal Components ($N_{\text{sys}}$) used to model the background noise (correlated between detectors) is a critical step in the analysis. This choice is guided by the goal of minimizing the final uncertainty on the signal amplitude ($\sigma_{T}$). Although the noise characteristics can differ slightly between the two datasets, leading to slightly different optimal PC counts, we chose a single number of components for consistency. As shown in Figure 6, we selected $N_{\text{sys}}=10$ for both mirror positions. This value represents a robust choice, located in the region where the signal uncertainty is at its minimum, ensuring a balanced and effective noise subtraction.
@@ 3.4 Final analysis
To fit our model to the data, we use a linear regression framework based on Maximum Likelihood Estimation (MLE). For each pixel, we construct a likelihood function based on the assumption of Gaussian noise.
By maximizing this function, we find the best-fit values for all the model parameters: the offset, the contribution of each Principal Component $T^{\mathrm{syst}}_{i}(t)$, and the amplitude of the signal $T_{\text{S}}$. This provides a robust noise model for each pixel, which can then be subtracted to search for a significant signal. For pixels near the expected trajectory, the signal component, proportional to the Gaussian function in Eq.(6), is included in the fit, allowing us to directly constrain its amplitude.
@@
Figure 7: Performance of the noise subtraction procedure. Left: Time evolution of the raw temperature data (solid line) compared to the model (dashed line) for a representative pixel, showing the model’s ability to accurately track large-scale thermal fluctuations. Center: Resulting temperature residuals for all pixels in the signal region after model subtraction; the fluctuations are centered around zero with a significantly reduced amplitude compared to the raw data. Right: Histogram of the temperature residuals from all pixels, fitted with a Gaussian profile to demonstrate that the remaining noise is stochastic.
@@ 4 Exclusion Curve
The next step is to translate this temperature limit into a limit on the power incident on the detector. The relationship between power and temperature variation $\Delta T$, for a blackbody detector, is described by Stefan-Boltzmann’s law, integrated over the bandwidth $H(\nu)$, $A\Omega$ is the throughput of each pixel, $\zeta(\nu)$ is the efficiency of the optical system composed of the three lenses (with $\langle\zeta(\nu)\rangle=0.6$), and $\langle\eta(\nu)\rangle=\langle\zeta(\nu)\cdot H(\nu)\rangle=0.3$ represents the detection efficiency seen in equation 4. The power amplitude is given by:
$A_{\text{power}}=A\Omega\int_{\nu_{min}}^{\nu_{max}}\zeta(\nu)H(\nu)\Delta B(\Delta T,\nu)d\nu.$ (7)
In our operating regime, where the total system temperature is high relative to the photon energy we want to measure ($k_{B}T_{\text{tot}}\gg h\nu$), we can use the Rayleigh-Jeans approximation, so we have $\Delta B=B(T_{background}+\Delta T,\nu)-B(T_{background},\nu)$. This condition is largely satisfied, especially for the 300 K lens, where the thermal energy is around 25 times greater than the energy of the 244 GHz photon. By calculating the integral, taking into account the measured bandwidth and the limit $\Delta T_{\text{lim}}$, we can convert the temperature limit into a detectable power limit. This limit on power is then converted into a limit on kinetic mixing coupling $\chi$ (Eq. 4). The final expression for the coupling is: